OpenStudy (anonymous):

Part 1. Using the two functions listed below, insert numbers in place of the letters a, b, c, and d so that f(x) and g(x) are inverses. f(x)= x+a over b g(x)=cx−d

OpenStudy (anonymous):

I'm so confused. Please explain and not just give me the answer!

OpenStudy (campbell_st):

well basically a and d have the same value as does b and c so pick some numbers

OpenStudy (anonymous):

a=2 and d=6 i guess

OpenStudy (campbell_st):

ok so \[f(x) = \frac{x + 2}{b}\] then d is also 2 \[g(x) = cx - 2\]

OpenStudy (campbell_st):

so b and c have the same value so b = 3 then c = 3 \[f(x) = \frac{x + 2}{3}~~~~g(x) = 3x - 2\]

OpenStudy (anonymous):

Oh okay.

OpenStudy (anonymous):

I also have to show work to prove that they are inverses. Would i just solve the equations?

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

ok... so look at the 1st equation \[y = \frac{c + a}{b}\] so find th inverse swap x and y, then make y the subject \[x = \frac{y + a}{b}\] multiply both sides by b \[bx = y + a\] subtract a \[bx - a = y~~~or~~~y = bx - a\] this is the inverse of f(x) now compare it to g(x) = cx - d o you can say b = c and a = d when you take the inverses